OpenStudy (hxkage):

If cosec A/ sec B = p and cot A/ sec B = q, prove that:

OpenStudy (hxkage):

\[(p^{2} - q^{2}) . \cot^{2}a= q ^{2}\]

zepdrix (zepdrix):

This one seems pretty straight forward, just plug everything in for the left side of the equation.

zepdrix (zepdrix):

\[\large\rm \left(\left[\frac{\csc A}{\sec B}\right]^2-\left[\frac{\cot A}{\sec B}\right]^2\right)\cot^2A\]I replaced my p and q. And then just simplify from there, ya?

OpenStudy (hxkage):

I see

OpenStudy (hxkage):

Using the identity, we get \[(\frac{ 1 }{ secB } )\cot ^{2}A\]

OpenStudy (hxkage):

Yea?

zepdrix (zepdrix):

Hmm the denominators should be squared as well, ya?\[\large\rm \left(\frac{\csc^2A}{\sec^2B}-\frac{\cot^2A}{\sec^2B}\right)\cot^2A\]

zepdrix (zepdrix):

\[\large\rm \left(\frac{\csc^2A-\cot^2A}{\sec^2B}\right)\cot^2A\]But yes, your Pythagorean Identity cleans up the numerator,\[\large\rm \left(\frac{1}{\sec^2B}\right)\cot^2A\]

OpenStudy (hxkage):

Dang, didn't notice that ;-;

zepdrix (zepdrix):

Haha I love how you called it "the" identity. We're in trig XD the land of identities. But I knew which one you were talking about :D lol

OpenStudy (hxkage):

XD I suck at trig btw ;-;

OpenStudy (hxkage):

So, it's always 'the' identity :P

OpenStudy (hxkage):

So now we get \[(\frac{ \cot ^{2}A }{ \sec ^{2} B }\]

OpenStudy (hxkage):

Forgot to close the bracket -.-

zepdrix (zepdrix):

k looks good.

OpenStudy (hxkage):

 Yea?